OpenStudy (anonymous):

Calculate the area of the triangle with the given vertices: A(2, 0, 0), B (0, 2, 0) and C(0, 0, 2) ?

OpenStudy (anonymous):

i figured out AB =(2,0,-2) AC = (0,2,-2) & BC = (-2,2,0)

OpenStudy (anonymous):

what to do from there?

OpenStudy (anonymous):

AB x AC You know vectors?

OpenStudy (anonymous):

<2,-2,0> <2,0,-2>

OpenStudy (anonymous):

why do i cross product them?

OpenStudy (anonymous):

It is a vector method of finding the area. What method are you attempting (1/2)bh

OpenStudy (anonymous):

i was thinking of finding the absolute value of AB, AC, and BC for the length of each side

OpenStudy (anonymous):

then doing something from there

OpenStudy (anonymous):

so i can cross product any of the of length coordinates and get an area?

OpenStudy (anonymous):

Cross sides AB and AC. And find the magnitude. That is the area of the related parallelogram; half that is the area of your triangle. The area is\[2 \sqrt{3}\]Try your method, see if you get the same thing.

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

didnt get same answer but i never knew cross product was used to find area...ty

OpenStudy (anonymous):

Let's see the work from your method.

OpenStudy (anonymous):

What is this? 'i figured out AB =(2,0,-2) AC = (0,2,-2) & BC = (-2,2,0)' Is this the vector AB? and AC?

OpenStudy (anonymous):

from there i got |AB|= |AC| = |BC| = 2sqrt(2)

OpenStudy (anonymous):

oh i get the right answer

OpenStudy (anonymous):

but had to assume it was a right angle triangle

OpenStudy (anonymous):

with all the lengths and just assuming its a right angle triangle i figure out height: b^2=c^2-a^2 and plug everything in

OpenStudy (anonymous):

but cross product is soo much faster

OpenStudy (anonymous):

Powerful tool.